OpenStudy (anonymous):

Find the value of the positive integer n for which division of x^(2n+1) +3x^2 - 28 by x-4 gives the remainder of 52.

OpenStudy (anonymous):

\[x ^{2n+1}+3x^2-28 = 52 \mod(7)\] i don't know if you can understand this

OpenStudy (anonymous):

Okay,but what is mathchoice mod (7) ?

OpenStudy (anonymous):

\[means : x^{2n+1}+3x^2-28=7k+52, k \in \mathbb{Z}^+\]

OpenStudy (anonymous):

\[x ^{2n+1}+3x^2−28≡52\mod(7) \] better i guess

OpenStudy (anonymous):

Okay,I'll try it first